Hello, I am having immense difficulty solving this problem:
(x-3)^2 = 1
I have tried multiplying out (x-3) times (x-3) which equals x^2 - 6x + 9, and then subtracting the one from both sides, but something is not working and I am still very confused. Please help?

Answer by Alan3354(69443)

Hello, I am having immense difficulty solving this problem:
(x-3)^2 = 1
----------
Get the sqrt of both sides:
x-3 = 1
x = 4
========
Also,
x-3 = -1
x = 2